# Adsorption of polyelectrolytes in the presence of varying dielectric discontinuity between solution and substrate

### v2

@@ III.1.2 Influence of multivalent counterions
The key to successful deposition of PE multilayers in a layer-by-layer fashion here would be charge inversion that should occur after each adsorption step. In $3:-1$ salt and absence of any substrates (dashed line in Fig. 3(b)), the overcompensation of DNA charge by trivalent cations condensed on the PE leads to the charge reversal of the PE and, consequently, the reversal of $\phi$ Yang et al. (2023); Vahid et al. (2024). However, the proximity of the substrate impedes full condensation of cations around the PE. This spatial limitation results in a diminished potential when the PE is near a non-polarizable substrate, as compared to the case where no substrate is present (dashed line). In addition, we observe a notable effect of different substrate types on the charge reversal of the PE. the potential undergoes rapid decrease as a function of $\varepsilon_{\rm s}$. The proximity of low dielectric constant substrates to the PE tends to enhance charge reversal, whereas high dielectric constant, metallic-character substrates appear to suppress it. This phenomenon can be explained by considering the nature of induced charges on the low dielectric constant substrates. These charges have the same sign as the PE, attracting more counterions to the vicinity of the substrate and facilitating charge reversal. In contrast, the induced charges on high dielectric constant, metallic-character substrates, being opposite in sign to those of the PE, repel nearby counterions, thereby suppressing charge reversal. This can be seen in the total ion charge profiles presented in Figs. 3(c)-(e). The density of positive charges around the PE close to the substrate with $\varepsilon_{\rm s}=2$ is higher than that with $\varepsilon_{\rm s}=8000$, which leads to charge reversal, and consequently the reversal of the electrostatic potential. Additionally, we have calculated the value of $\phi$ in cases when enough space exists for a layer of counterions between the PE and the substrate (see Fig. S4 of SM). We observe that under these conditions, the results for different values of $\varepsilon_{\rm s}$ converge to those observed for non-polarizable substrates.
Figure 4: Normalized potential of mean force ($\beta V/L$) versus distance between the center of the PE and the substrate plane in $1:-1$ (first row), $2:-1$ (second row), and $3:-1$ (third row) salt solutions, across varying salt concentrations $Z^{\rm c}c^{\rm c}$. Here, $\varepsilon_{\rm s}=8000$. The vertical dashed lines represent the distance at which the surface of the PE is in contact with the substrate beads $z_{1}=(\sigma^{\rm PE}+\sigma^{\rm s})/2$, while the dotted lines represent the distance at which they are one ion diameter apart such that $z_{2}=z_{1}+\sigma^{i}$.
III.1.3 Influence of PE line charge density
@@
As already discussed regarding strongly charged PEs ($\lambda_{0}=-6$ $e/$nm) in Fig. 2, the presence of multivalent cations in salt leads to more effective screening of induced charges as well as the PE charge, thereby weakening the electrostatic attraction towards the substrate. In the case of $3:-1$ salt, the highly charged PE is completely repelled from the substrate. Interestingly, multivalent cations do not significantly influence the interaction of weakly charged PEs ($\lambda_{0}=-0.3$ $e/$nm) with the substrate, and adsorption is possible with higher salt concentrations.
@@ III.2 Two PEs parallel to a high dielectric constant substrate
Figure 5: (a) A snapshot of the simulation box with two parallel identical PEs whose surfaces are separated by $D$ (their center-to-center separation is $D+\sigma^{\rm PE}$), measuring $20\times 20\times 30$ nm3 with periodic boundary conditions along the $x$ and $y$ axes. The substrate, depicted in yellow and with the dielectric constant of $\varepsilon_{\rm s}=8000$, is positioned at $z=0$, while a hard reflecting wall is located at $z=15$ nm. The system also contains monovalent counterions (red) and monovalent coions (blue) at $Z^{\rm c}c^{\rm c}=|Z^{\rm a}|c^{\rm a}=60$ mM. (b) Close-up view of the PEs oriented along the $y$ axis and the bead-based substrate which is along the $xy$ plane at $z=0$ and structured in a triangular lattice. $D$ is the PE-surface-to-PE-surface distance. (c)-(e) Panels depicting the total two-dimensional charge distribution $n_{\rm t}=n^{\rm c}+n^{\rm a}$ for cations and anions in $Z^{\rm c}:-1$ salt at $Z^{\rm c}c^{\rm c}=|Z^{\rm a}|c^{\rm a}=500$ mM. The data are averaged along the $y$ axis. The PEs are at $z=1.4$ nm along the $y$ axis, with an equilibrium spacing of $D=0.6$ nm between them. The PEs are shown in gray.
@@ III.2.1 PE-PE interactions on a high dielectric constant substrate
Figure 6: Normalized potential of mean force ($\beta V/L$) between PEs plotted against their surface-to-surface distance $D$. Data are presented for varying salt valency ratios ($Z^{\rm c}:Z^{\rm a}$, left to right), salt concentrations ($Z^{\rm c}c^{\rm c}$), and PE line charge $\lambda_{0}$. The substrate relative dielectric constant is $\varepsilon_{\rm s}=8000$. Both PE centers are fixed at $z=1.4$ nm from the substrate plane. Vertical dotted lines indicate $D=\sigma^{i}=0.5$ nm.
In Fig. 6, we present the normalized potential of mean force $\beta V/L$ as a function of the surface-to-surface distance $D$ between the PEs at $z=1.4$ nm. Here, we systematically vary both the valency of cations in the added salt and its concentration. We observe that the strongly charged PEs ($\lambda_{0}=-6$ $e/$nm) repel each other at all distances in $1:-1$ salt solutions, and $V$ approaches zero at all concentrations when $D>2$ nm. At high enough salt concentrations, when $Z^{\rm c}c^{\rm c}\geq 250$ mM, PEs attract each other in the presence of $2:-1$ salt, see panel (e). This attractive force becomes even more pronounced in $3:-1$ salt, where attraction appears at even lower concentrations (see panel (f)). This attraction is attributed to the bridging effect of the cations situated between the PEs and the correlation effects induced by these ions. Additionally, the results of Fig. S5 of the SM show that altering the substrate dielectric constant $\varepsilon_{\rm s}$ from $500$ to $8000$ has no significant impact on the interactions between PEs near the high dielectric constant, metallic-character substrates, particularly at higher salt concentrations.
@@
Figure 7: The normalized potential of mean force $\beta V/L$ as a function of the distance $z$ from the center of each PE to the substrate plane. The data are presented for salts with valencies $1:-1$, $2:-1$, and $3:-1$ while varying the salt concentration $Z^{\rm c}c^{\rm c}$ and PE line charge $\lambda_{0}$. The substrate relative dielectric constant is $\varepsilon_{\rm s}=8000$. The PEs are at their equilibrium distances of $D=0.25$ nm for $\lambda_{0}=-0.3$ $e/$nm and $D=0.55$ nm for $\lambda_{0}=-6$ $e/$nm from each other. Vertical dashed lines indicate where the PE surface contacts the substrate beads at $z_{1}=(\sigma^{\rm PE}+\sigma^{\rm s})/2$, and dotted lines show where the separation equals one ion diameter at $z_{2}=z_{1}+\sigma^{i}$.
@@ III.2.2 Interaction of a PE pair with a high dielectric constant substrate
The relatively strong attraction induced by charge correlations due to multivalent counterions leads to bridged PE chains together in the bulk of the solution. On the other hand, strongly charged PEs experience screening-reduced adsorption with increasing salt concentration, as shown in Fig. 4. The remaining question then concerns the influence of the second PE on the PE-substrate interactions. To obtain a dense adsorbate layer, an attractive interaction towards the substrate and a potential energy minimum should prevail upon adsorption on the substrate. To this end, we consider two PEs approaching the substrate at their corresponding minimum energy distances from each other at $D=0.25$ nm and $D=0.55$ nm for PE charge densities of $\lambda_{0}=-0.3$ $e/$nm and $\lambda_{0}=-6$ $e/$nm, respectively. In cases where PEs do not exhibit mutual attraction, such as for $\lambda_{0}=-6$ $e/$nm in $1:-1$ salt, we use the same $D$ value used for other salts at that particular $\lambda_{0}$.
@@
Figure 7 shows that the strongly charged PEs ($\lambda=-6$ $e/$nm) are repelled from the substrate at all salt concentrations here. The competition between electrostatic attraction and steric repulsion governs the adsorption process here. Initially, condensation of cations between the two PEs (see Fig. 5(c)-(e)) results in the neutralization of charges induced on the substrate, thereby diminishing the electrostatic attraction towards the substrate. In addition, the steric repulsion, arising from the condensed ions between the substrate and the PEs, becomes the dominant force, leading to repulsion of the PEs. We note that a higher salt concentration enhances the ion condensation on the PEs, thereby augmenting this repulsive effect. On the other hand, PEs with a low charge density $\lambda=-0.3$ $e/$nm can be adsorbed to the substrate even for monovalent salt, and the adsorption strength increases as the salt concentration increases. In this case, similar to the single PE case shown in Fig. 4, there is no energy barrier impeding the adsorption of the PEs to the substrate due to the reduced attraction of cations to the PEs. Furthermore, the entropic depletion effect leads to the PEs being attracted to the substrate, an effect that enhances with the rise in salt concentration. Similar findings were reported in Ref. Harmat et al., 2023, where interactions of poly-L-lysine and poly-L-arginine with $\alpha$-quartz surfaces were studied, with sodium ions as counterions. It was observed that at low pH levels, where the peptides are strongly charged, significant electrostatic repulsion between peptides leads to reduced substrate coverage. Conversely, at higher pH levels, the reduced charge on the peptides enhances the formation of denser peptide films. Despite increased substrate coverage, these films exhibit decreased stability. Our results in Fig. 7 further show that increasing cation valency to $Z^{\rm c}=3$ does not notably affect the interaction between the PEs and the substrate.
@@
Figure 8: The normalized binding energy of the PEs to the substrate $\beta E_{\rm B}/L$ as a function of PE line charge $\lambda_{0}$ and surface-to-surface PE distance $D$. The data correspond to $1:-1$ salt at concentration $Z^{\rm c}c^{\rm c}=60$ mM (left) and $Z^{\rm c}c^{\rm c}=500$ mM (right). The substrate relative dielectric constant is $\varepsilon_{\rm s}=8000$. Simulated values are marked with symbols; circles represent cases where the PEs attract each other, and crosses indicate cases where the PEs repel each other. The color scheme is based on 2D interpolation between the discrete measurements. The maximum value of the color bar is capped at $0.6$ nm-1. The higher values are shown in dark blue.
@@
To summarize the findings of this section, we calculate the binding energy $E_{\rm B}$, which is defined as the energy required to bring a bound PE pair from their equilibrium distance close to the substrate to infinity and can be quantified by twice the absolute value of the (negative) minimum of $V$ in Fig. 7, where $V_{\rm min}=V(z_{\rm min})$, with $z_{\rm min}\approx 1.4$ nm. This distance is the equilibrium distance for cases where adsorption to the substrate is observed. For consistency, the same value of $z_{\rm min}$ is also used for cases where adsorption is not observed. In Fig. 8, we explore the dependence of $E_{\rm B}$ on both $D$ and $\lambda_{0}$ in $1:-1$ salt solution at two concentrations in the case of a high dielectric constant, metallic-character substrate with $\varepsilon_{\rm s}=8000$. This case has been chosen since, from the data presented above, it is the most promising setup for the production of a densely populated monolayer. At the lowest salt concentration $Z^{\rm c}c^{\rm c}=60$ mM (left), the PE-substrate interaction is attractive when $|\lambda_{0}|<3$ $e/$nm, and the PEs are sufficiently close, i.e., $D<0.45$ $e/$nm. However, the PEs themselves do not attract each other (cf. Fig. 6(a)). At the highest salt concentration $Z^{\rm c}c^{\rm c}=500$ mM (right), the PEs are attracted to the substrate when $|\lambda_{0}|<1.5$ $e/$nm for all $D$. On the other hand, PE-PE attraction is observed when $D<0.45$ nm and $|\lambda_{0}|<0.6$ $e/$nm (cf. Fig. 6(a) and Fig. S7 of the SM), denoted with circles in Fig. 8. These relatively few points are the only cases for which the formation of a dense monolayer can be expected on a high dielectric constant, metallic-character neutral substrate.

### v3

@@ III.1.2 Influence of multivalent counterions
The key to successful deposition of PE multilayers in a layer-by-layer fashion here would be charge inversion that should occur after each adsorption step. In $3:1$ salt and absence of any substrates (dashed line in Fig. 3(b)), the overcompensation of DNA charge by trivalent cations condensed on the PE leads to the charge reversal of the PE and, consequently, the reversal of $\phi$ Yang et al. (2023); Vahid et al. (2024). However, the proximity of the substrate impedes full condensation of cations around the PE. This spatial limitation results in a diminished potential when the PE is near a non-polarizable substrate, as compared to the case where no substrate is present (dashed line). In addition, we observe a notable effect of different substrate types on the charge reversal of the PE. the potential undergoes rapid decrease as a function of $\varepsilon_{\rm s}$. The proximity of low dielectric constant substrates to the PE tends to enhance charge reversal, whereas high dielectric constant, metallic-character substrates appear to suppress it. This phenomenon can be explained by considering the nature of induced charges on the low dielectric constant substrates. These charges have the same sign as the PE, attracting more counterions to the vicinity of the substrate and facilitating charge reversal. In contrast, the induced charges on high dielectric constant, metallic-character substrates, being opposite in sign to those of the PE, repel nearby counterions, thereby suppressing charge reversal. This can be seen in the total ion charge profiles presented in Figs. 3(c)-(e). The density of positive charges around the PE close to the substrate with $\varepsilon_{\rm s}=2$ is higher than that with $\varepsilon_{\rm s}=8000$, which leads to charge reversal, and consequently the reversal of the electrostatic potential. Additionally, we have calculated the value of $\phi$ in cases when enough space exists for a layer of counterions between the PE and the substrate (see Fig. S4 of SM). We observe that under these conditions, the results for different values of $\varepsilon_{\rm s}$ converge to those observed for non-polarizable substrates.
III.1.3 Influence of PE linear charge density
@@
Figure 4: Normalized potential of mean force ($\beta V/L$) versus distance between the center of the PE and the substrate plane in (a)-(b) $1:1$, (c)-(d) $2:1$, (e)-(f) and $3:1$ salt solutions, across varying salt concentrations $Z^{\rm c}c^{\rm c}$. Here, $\varepsilon_{\rm s}=8000$. The vertical dashed lines represent the distance at which the surface of the PE is in contact with the substrate beads $z_{1}=(\sigma^{\rm PE}+\sigma^{\rm s})/2$, while the dotted lines represent the distance at which they are one ion diameter apart such that $z_{2}=z_{1}+\sigma^{i}$.
As already discussed regarding strongly charged PEs ($\lambda_{0}=-6$ $e/$nm) in Fig. 2, the presence of multivalent cations in salt leads to more effective screening of induced charges as well as the PE charge, thereby weakening the electrostatic attraction towards the substrate. In the case of $3:1$ salt, the highly charged PE is completely repelled from the substrate. Interestingly, multivalent cations do not significantly influence the interaction of weakly charged PEs ($\lambda_{0}=-0.3$ $e/$nm) with the substrate, and adsorption is possible with higher salt concentrations.
@@ III.2 Two PEs parallel to a high dielectric constant substrate
Figure 5: (a) A snapshot of the simulation box with two parallel identical PEs whose surfaces are separated by $D$ (their center-to-center separation is $D+\sigma^{\rm PE}$), measuring $20\times 20\times 30$ nm3 with periodic boundary conditions along the $x$ and $y$ axes. The substrate, depicted in yellow and with the dielectric constant of $\varepsilon_{\rm s}=8000$, is positioned at $z=0$, while a hard reflecting wall is located at $z=15$ nm. The system also contains monovalent counterions (red) and monovalent coions (blue) at $Z^{\rm c}c^{\rm c}=|Z^{\rm a}|c^{\rm a}=60$ mM. (b) Close-up view of the PEs oriented along the $y$ axis and the bead-based substrate which is along the $xy$ plane at $z=0$ and structured in a triangular lattice. $D$ is the PE-surface-to-PE-surface distance. (c)-(e) Panels depicting the total two-dimensional charge distribution $n_{\rm t}=n^{\rm c}+n^{\rm a}$ for cations and anions in $Z^{\rm c}:1$ salt at $Z^{\rm c}c^{\rm c}=|Z^{\rm a}|c^{\rm a}=500$ mM. The data are averaged along the $y$ axis. The PEs are at $z=1.4$ nm along the $y$ axis, with an equilibrium spacing of $D=0.6$ nm between them. The PEs are shown in gray.
@@ III.2.1 PE-PE interactions on a high dielectric constant substrate
Figure 6: Normalized potential of mean force ($\beta V/L$) between PEs plotted against their surface-to-surface distance $D$. Data are presented for varying salt valency ratios ($Z^{\rm c}:|Z^{\rm a}|$, left to right), salt concentrations ($Z^{\rm c}c^{\rm c}$), and PE linear charge density of (a)-(c) $\lambda_{0}=-0.3$ $e/$nm and (d)-(f) $\lambda_{0}=-6$ $e/$nm. The substrate relative dielectric constant is $\varepsilon_{\rm s}=8000$. Both PE centers are fixed at $z=1.4$ nm from the substrate plane. Vertical dotted lines indicate $D=\sigma^{i}=0.5$ nm.
In Fig. 6, we present the normalized potential of mean force $\beta V/L$ as a function of the surface-to-surface distance $D$ between the PEs at $z=1.4$ nm. Here, we systematically vary both the valency of cations in the added salt and its concentration. We observe that the strongly charged PEs ($\lambda_{0}=-6$ $e/$nm) repel each other at all distances in $1:1$ salt solutions, and $V$ approaches zero at all concentrations when $D>2$ nm. At high enough salt concentrations, when $Z^{\rm c}c^{\rm c}\geq 250$ mM, PEs attract each other in the presence of $2:1$ salt, see panel (e). This attractive force becomes even more pronounced in $3:1$ salt, where attraction appears at even lower concentrations (see panel (f)). This attraction is attributed to the bridging effect of the cations situated between the PEs and the correlation effects induced by these ions. Additionally, the results of Fig. S5 of the SM show that altering the substrate dielectric constant $\varepsilon_{\rm s}$ from $500$ to $8000$ has no significant impact on the interactions between PEs near the high dielectric constant, metallic-character substrates, particularly at higher salt concentrations.
@@
Figure 7: The normalized potential of mean force $\beta V/L$ as a function of the distance $z$ from the center of each PE to the substrate plane. The data are presented for salts with valencies $1:1$, $2:1$, and $3:1$ while varying the salt concentration $Z^{\rm c}c^{\rm c}$. The PE linear charge density is (a)-(c) $\lambda_{0}=-0.3$ $e/$nm and (d)-(f) $\lambda_{0}=-6$ $e/$nm. The substrate relative dielectric constant is $\varepsilon_{\rm s}=8000$. The PEs are at their equilibrium distances of $D=0.25$ nm for $\lambda_{0}=-0.3$ $e/$nm and $D=0.55$ nm for $\lambda_{0}=-6$ $e/$nm from each other. Vertical dashed lines indicate where the PE surface contacts the substrate beads at $z_{1}=(\sigma^{\rm PE}+\sigma^{\rm s})/2$, and dotted lines show where the separation equals one ion diameter at $z_{2}=z_{1}+\sigma^{i}$.
@@ III.2.2 Interaction of a PE pair with a high dielectric constant substrate
The relatively strong attraction induced by charge correlations due to multivalent counterions leads to bridged PE chains together in the bulk of the solution. On the other hand, strongly charged PEs experience screening-reduced adsorption with increasing salt concentration, as shown in Fig. 4. The remaining question then concerns the influence of the second PE on the PE-substrate interactions. To obtain a dense adsorbate layer, an attractive interaction towards the substrate and a potential energy minimum should prevail upon adsorption on the substrate. To this end, we consider two PEs approaching the substrate at their corresponding minimum energy distances from each other at $D=0.25$ nm and $D=0.55$ nm for PE charge densities of $\lambda_{0}=-0.3$ $e/$nm and $\lambda_{0}=-6$ $e/$nm, respectively. In cases where PEs do not exhibit mutual attraction, such as for $\lambda_{0}=-6$ $e/$nm in $1:1$ salt, we use the same $D$ value used for other salts at that particular $\lambda_{0}$.
@@
Figure 7 shows that the strongly charged PEs ($\lambda=-6$ $e/$nm) are repelled from the substrate at all salt concentrations here. The competition between electrostatic attraction and steric repulsion governs the adsorption process here. Initially, condensation of cations between the two PEs (see Fig. 5(c)-(e)) results in the neutralization of charges induced on the substrate, thereby diminishing the electrostatic attraction towards the substrate. Increasing the cation valency lowers the steric barrier height at every salt concentration, possibly due to the reduced number of ions in the system. In addition, the steric repulsion, arising from the condensed ions between the substrate and the PEs, becomes the dominant force, leading to repulsion of the PEs. This aligns with theoretical findings using PB theory, which have shown that DNA-cationic lipid complexes with monovalent counterions become unstable when additional DNA strands are adsorbed onto the substrate Bruinsma (1998); Cherstvy (2007). We note that a higher salt concentration enhances the ion condensation on the PEs, thereby augmenting this repulsive effect. On the other hand, PEs with a low charge density $\lambda=-0.3$ $e/$nm can be adsorbed to the substrate even for monovalent salt, and the adsorption strength increases as the salt concentration increases. In this case, similar to the single PE case shown in Fig. 4, there is no energy barrier impeding the adsorption of the PEs to the substrate due to the reduced attraction of cations to the PEs. Furthermore, the entropic depletion effect leads to the PEs being attracted to the substrate, an effect that enhances with the rise in salt concentration. Similar findings were reported in Ref. Harmat et al., 2023, where interactions of poly-L-lysine and poly-L-arginine with $\alpha$-quartz surfaces were studied, with sodium ions as counterions. It was observed that at low pH levels, where the peptides are strongly charged, significant electrostatic repulsion between peptides leads to reduced substrate coverage. Conversely, at higher pH levels, the reduced charge on the peptides enhances the formation of denser peptide films. Despite increased substrate coverage, these films exhibit decreased stability. Our results in Fig. 7 further show that increasing cation valency to $Z^{\rm c}=3$ does not notably affect the interaction between the PEs and the substrate.
@@
Figure 8: The normalized binding energy of the PEs to the substrate $\beta E_{\rm B}/L$ as a function of PE linear charge density $\lambda_{0}$ and surface-to-surface PE distance $D$. The data correspond to $1:1$ salt at concentration $Z^{\rm c}c^{\rm c}=60$ mM (left) and $Z^{\rm c}c^{\rm c}=500$ mM (right). The substrate relative dielectric constant is $\varepsilon_{\rm s}=8000$. Simulated values are marked with symbols; circles represent cases where the PEs attract each other, and crosses indicate cases where the PEs repel each other. The color scheme is based on 2D interpolation between the discrete measurements. The maximum value of the color bar is capped at $0.6$ nm-1. The higher values are shown in dark blue. Figure S8 of the SM presents $\beta V/L$ between PEs and the substrate as a function of their distance $z$ for selected points from this figure.
@@
To summarize the findings of this section, we calculate the binding energy $E_{\rm B}$, which is defined as the energy required to bring a bound PE pair from their equilibrium distance close to the substrate to infinity and can be quantified by twice the value of the (negative) minimum of $V$ in Fig. 7, where $V_{\rm min}=V(z_{\rm min})$, with $z_{\rm min}\approx 1.4$ nm. This distance is the equilibrium distance for cases where adsorption to the substrate is observed. For consistency, the same value of $z_{\rm min}$ is also used for cases where adsorption is not observed. In Fig. 8, we explore the dependence of $E_{\rm B}$ on both $D$ and $\lambda_{0}$ in $1:1$ salt solution at two concentrations in the case of a high dielectric constant, metallic-character substrate with $\varepsilon_{\rm s}=8000$. This case has been chosen since, from the data presented above, it is the most promising setup for the production of a densely populated monolayer. At the lowest salt concentration $Z^{\rm c}c^{\rm c}=60$ mM (left), the PE-substrate interaction is attractive when $|\lambda_{0}|<3$ $e/$nm, and the PEs are sufficiently close, i.e., $D<0.45$ nm. However, the PEs themselves do not attract each other (cf. Fig. 6(a)). At the highest salt concentration $Z^{\rm c}c^{\rm c}=500$ mM (right), the PEs are attracted to the substrate when $|\lambda_{0}|<1.5$ $e/$nm for all $D$. On the other hand, PE-PE attraction is observed when $D<0.45$ nm and $|\lambda_{0}|<0.6$ $e/$nm (cf. Fig. 6(a) and Fig. S7 of the SM), denoted with circles in Fig. 8. These relatively few points are the only cases for which the formation of a dense monolayer can be expected on a high dielectric constant, metallic-character neutral substrate.